find the value for m and b in the following system so that the solution to the sysytem is (-3, 4)
5x+7y =b
and
mx+y = 22
thank you for the help

Answer by uma(370)

5x+7y =b
mx+y = 22
Given that (-3,4) is a solution to the above system
Plugging in x = -3 and y = 4 in the first equation,
5(-3)+7(4) = b
==> -15+28 = b
==> 13 = b
So b = 13
Now m(-3) + 4 = 22
==> -3m + 4 = 22
==> -3m = 22-4
==> -3m = 18
==> m = 18/-3
==> m = -6
Good Luck!!!